OpenStudy (anonymous):

y = 2x + 6 4x − 3y = −6 what is x and y

OpenStudy (gw2011):

Substitute y=2x+6 into the equation 4x-3y=-6. 4x-3(2x+6)=-6 4x-6x-18=-6 -2x=-6+18 -2x=12 x=-6 Substitute x=-6 into y=2x+6 to find y y=2x+6 y=2(-6)+6 y=-12+6 y=-6 Therefore, x=-6 and y=-6